OpenStudy (celinegirl):

When the function f(x) = 3(5)x is changed to f(x) = 3(5)x + 22, what is the effect? There will be no change to the graph because the exponential portion of the function remains the same. The y-intercept is 22 spaces higher. The x-intercept is 22 spaces higher. All input values are moved 22 spaces to the right.

OpenStudy (aihberkhan):

Well we know that the +22 is the y-value. Now "+" means more, higher, etc. So now what can you conclude? @celinegirl

OpenStudy (celinegirl):

B?

OpenStudy (idku):

y=F(x)+C is shifted C units up (if C is a positive number) from the parent function y=F(x)

OpenStudy (idku):

And if you had y=F(x)-C (where C is positive, and thereby -C is a negative number) then that is shifted C units down, from the parent function y=F(x)

OpenStudy (idku):

y=F(x+C) shift left by C units from y=F(x). y=F(x-C) shift right by C units from y=F(x). (where C, again, is a positive number/constant)

OpenStudy (idku):

y=C•F(x) For positive C, such that 0<C<1 you are stretching the function When C=1, nothing happens to y=F(x). For positive C, such that C>1 you are shrinking the function (taller but thinner)

OpenStudy (idku):

y=C•F(x) For negative C, such that 0<C<1 you are stretching the function, and reflecting the function over the x-axis. When C=-1, just reflects the function over the x-axis. For negative C, such that C>1 you are shrinking the function (taller but thinner), and reflecting the function over the x-axis.

OpenStudy (idku):

y=F(-x) is a reflection over the y-axis. (Sometimes reflection over the y-axis and the x-axis can be the same action, and that is when the F(x) is a line)

OpenStudy (celinegirl):

@idku hm I think I got some of that ha

OpenStudy (pawanyadav):

It's not a single time ,it should be twice ha ha